# Personalized Path Recourse for Reinforcement Learning Agents

### v2

@@ 4.1 Reward Shaping
Similarity
Measuring the similarity between two sequences, $\tau_{0}$ and $\tau_{r}$, is a non-trivial task. The problem becomes even more challenging if the sequences have different lengths. For this reason, traditional metrics such as Euclidean or Manhattan distance metrics do not apply. In our work, we measure the distance between $\tau_{0}$ and any route $\tau_{r}$ by Levenshtein distance (Levenshtein et al., 1966), denoted as $d(\tau_{0},\tau_{r})$, which counts the minimum number of operations (insertion, deletion, and substitution ) to convert $\tau_{0}$ to $\tau_{r}$. Then the similarity between $\tau_{0}$ and $\tau_{r}$ is computed by
@@
$f_{\text{sim}}(\cdot)$ value is between 0 and 1, where $f_{\text{sim}}(\tau_{0},\tau_{r})=1$ if $\tau_{0}$ and $\tau_{r}$ are identical, and close to 0 if they are very different. $f_{\text{sim}}(\cdot)$ can be used as the similarity reward $R_{\text{sim}}$, which is defined at the sequence level.
Personalization
We want the path to be personalized based on the behavior policy $\mathcal{P}_{A}$. This can be achieved by including a personalization reward, $R_{\text{pe}}$, in the reward function. We construct such a reward function from agent $A$’s policy $\mathcal{P}_{A}$, which is a probability distribution over states and action pairs. Thus, given a state and action pair $(\mathbf{s},a)$, $\mathcal{P}_{A}$ returns a probability, which represents how likely agent $A$ will take action $a$ for a given state $\mathbf{s}$. We want to use this value to represent whether a recommended action “respects” agent $A$’s behavior, i.e., the distribution. Thus, we define the personalization reward for each state-action pair as:
$R_{\text{pe}}(\mathbf{s},a)=h(\mathcal{P}_{A}(\mathbf{s},a)).$ (3)
Directly using the probability as the reward value (i.e., $h(p)=p$) could be misleading because the magnitude of the probabilities is a result of the cardinality of the action space. For example, if $|\mathcal{A}|=2$, a probability of 0.3 is considered low, and the personalization reward should also be low. However, when $|\mathcal{A}|$ is a very large value, 0.3 should be considered large, and the corresponding personalization reward should also be large. Thus, a linear relationship between the reward and probability is insufficient. Therefore, we design a link function $h(\cdot)$ that needs to satisfy the conditions as follows:
@@
Then, the total personalization reward defined at the sequence level is the sum of rewards for each state-action pair along the path:
$R_{\text{pe}}(\tau_{r})=\sum_{\mathbf{s}_{t}\in\tau_{r}}R_{\text{pe}}(\mathbf{s}_{t},a_{t}).$ (5)
@@
$R(\tau_{r})=R_{\text{goal}}(\tau_{r})+\lambda_{\text{sim}}R_{\text{sim}}(\tau_{r},\tau_{0})+\lambda_{\text{pe}}R_{\text{pe}}(\tau_{r}).$ (6)
@@ 4.2 Training
We use DQN to train a recourse agent $\mathcal{\pi}_{A}^{r}$ that can generate a recourse path $\tau_{r}$ for $\tau_{0}$ executed by agent $\mathcal{P}_{A}$, such that $\tau_{r}$ maximizes the reward $R(\tau_{r})$. We design an exploration strategy to help the training process converge faster.
@@
We want to encourage the agent to explore states and actions with high state-action values, avoid repetition of the same states/actions, and pay higher priority to the new states/actions. We achieve those goals by using an idea that is similar to Upper Confidence Bound (UCB) bandit algorithm. We define an exploration score for taking action $a_{t}$ from a state $\mathbf{s}_{t}$ at time $t$ as follows:
@@
The training algorithm is presented in Algorithm 1 to train $\pi_{A}^{r}$. First, we explore the environment using the strategy in equation 7 and record all the states, actions as well as corresponding rewards (lines 1-9). Next, we add the $k$ best records to buffer replay $B$ to train network $Q_{\theta}$, and update network $Q_{\theta}^{\prime}$ every $C$ step in the while loop (lines 10-16). Lines 12 converts the reward from trajectory-level to state-action level discounted return to update the network using Equation 1. This process is repeated until it converges, i.e., the agent learns to achieve optimal reward.
Input: Original route $\tau_{0}$, personalized policy $\mathcal{P}_{A}$, an environment parameterized by state set $\mathcal{S}$, action set $\mathcal{A}$, and reward function $R(.)$
@@
1 Initialize a replay buffer $B$ size $b$ Initialize two networks $Q_{\theta},Q_{\theta^{\prime}}$ with random parameters $\theta$ and $\theta^{\prime}$ while not converge do
2       $B=\emptyset$ for i=1 to $b$ do
3             Sample a path $\tau=<\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{T}>$ using equation 7 with decay $\epsilon$-greedy strategy Record all the action $a_{t}$ in the path Compute the reward $R(\tau)$ with equation 6
@@ 2       $B=\emptyset$ for i=1 to $b$ do
5      Collect $k$ best paths with the highest $R(\tau)$. foreach $k$ record do
@@ 8      Samples tuples from $B$ and update $\theta$ for $Q_{\theta}$:
($\alpha$ is the learning rate and $L$ is the loss function 1) Set $\theta^{\prime}=\theta$ every $C$ steps in the while loop
@@ 4.3 PPR for Supervised Learning
In supervised learning, $\tau$ is not a sequence of states but a sequence data $\mathbf{x}$ with a label $y=Y_{0}$. A personalized recourse means to generate a new sequence $\mathbf{x}^{\prime}$ with a different label $Y_{1}$ (goal), while being similar to $\mathbf{x}$ (similarity) and being in-distribution (personalization). The goal reward can be defined as the probability that $\mathbf{x}^{\prime}$ is classified as label $Y_{1}$ by a pre-trained classifier. The similarity reward is defined the same way as in Section 4.1, which is the Levenshtein distance. However, other distance metrics such as Euclidean distance and Manhattan distance can also be used if the sequences have fixed lengths. To incorporate personalization reward, we train a generator model from the training data to represent agent $A$. For example, if it is text data, $\mathcal{P}_{A}$ is a language model; if it is a sequence of vectors, $\mathcal{P}_{A}$ can be a simple RNN model that is either pre-trained or provided. Then the personalization policy is defined as the sampling probability of the next symbol, given the current prefix sequence: $\mathcal{P}_{A}(\mathbf{s_{t}},a_{t})=p(\mathbf{u_{t}}|\mathbf{u_{1}},...,\mathbf{u}_{t-1})$, where $\mathbf{u}_{i}$ represents each symbol in sequence data.
@@ 5 Experiments
In this section, we conduct experimental evaluations of PPR in three settings. In each setting, we generate agents with different behavior and then evaluate how PPR generates different recourse paths to adapt to those agents. The first two settings are reinforcement learning environments and the third one is text generation. Due to the space limit, we include the experiment in the supervised learning setting as well as other performance evaluation in the supplementary material.
@@
There are no baselines that work for all settings. We compare PPR against two methods, denoted as BL1 (Tsirtsis et al., 2021) and BL2 (Delaney et al., 2021), which work for some settings but not all. BL1 uses dynamic programming to exhaustively search all states and actions. Thus, it is not suitable for applications with very large state and action spaces, such as texts and atari games. BL2 is designed only for classification tasks in supervised learning settings and requires access to the training data for the classifier. Therefore, it is not suitable for reinforcement learning applications or text generation. The implementation procedures can be found in the supplementary material.
@@
We evaluated the generated paths in three aspects: personalization, goal satisfaction, and similarity, by reporting 3 types of scores $s_{\text{pe}},s_{\text{sim}},$ and $s_{\text{goal}}$, respectively. The goal satisfaction score $s_{\text{goal}}$ and similarity score $s_{\text{sim}}$ are the same as the goal reward and the similarity reward, i.e., $s_{\text{goal}}=R_{\text{goal}}$ and $s_{\text{sim}}=f_{\text{sim}}$ from equation 2. For personalization score $s_{\text{pe}}$, we report the average log of probabilities of the generated PPR if it is sampled from the personalized policy $\mathcal{P}_{A}$ the higher this value, the higher probability this PPR being generated from the personalized policy, i.e., the more this path follows the agent’s behavior. Thus, for a path $\tau_{r}=<\mathbf{s}_{1}^{r},...,\mathbf{s}_{T}^{r}>$, $s_{\text{pe}}=\frac{1}{T}\sum_{t=1}^{T}\log\>\mathcal{P}_{A}(\mathbf{s}^{r}_{t+1}|\mathbf{s}^{r}_{t},a^{r}_{t})$.
@@ 5.1 Grid-world environment
In this section, we apply PPR to a grid-world environment, shown in Figure 1. In this environment, a taxi driver wants to get to the destination (the white flag) as soon as possible. If the driver reaches the destination, the reward is 80. An extra reward of 30 will be given if the driver can collect money at the dollar sign. The reward for any other step is -1. Originally, the driver picked a bad route with a low reward, which either didn’t collect the money or took a detour to reach the destination. We randomly generate 10 such bad routes and an example is shown in Figure 1. We want to generate a better route that can help the taxi driver get a better reward, which means she will reach the destination faster and collect the money. More importantly, we want to be able to provide personalized recourse path that takes into consideration her driving habit. Also, the new route should not be too different from the original one - the similarity consideration.
Figure 1: The original route $\tau_{0}$ (the driver didn’t collect reward), the recourse paths for the experienced driver $\tau_{\text{exp}}$ and the recourse path for the new driver $\tau_{\text{new}}$.
We simulate two types of taxi drivers as two agents. Driver $A_{\text{exp}}$ is an experienced driver who’s comfortable driving on the highway, and driver $A_{\text{new}}$ is a newbie driver who prefers to avoid the highway and take local roads instead. We simulate these two drivers by designing different reward functions. Agent $A_{\text{exp}}$ is trained by setting higher rewards for the highway area and lower rewards for the local area and vice versa for training agent $A_{\text{new}}$333see the supplementary material for more detailed information.. Then the policy functions for the agents are provided as input to PPR.
We then run Algorithm 1 to generate personalized better routes for each agent. We set $\lambda_{\text{pe}}=0.1,\lambda_{\text{sim}}=0.1$ in Algorithm 1 (see the supplementary material for a sensitivity analysis of $\lambda_{\text{pe}}$ and $\lambda_{\text{sim}}$.) We show an example in Figure 1. As we can see, the recommended path $\tau_{\text{exp}}$ that is personalized for the experienced driver passes through the highway area because it follows the experienced driver policy. Likewise, the recommended route $\tau_{\text{new}}$ for the newbie driver takes a local road, also following his preferences reflected by the policy function. In this example, those new routes are the recourse paths of the original route, as they respect the driver’s driving behavior and facilitate a better outcome for the taxi while being similar to the old route from the starting point to the destination. Please refer to the supplementary material for all 10 routes and recourse paths.
@@
(a) $s_{\text{pe}}$, $s_{\text{sim}}$ ,$s_{\text{goal}}$ for PPR and BL1 trained on grid-world
(b) $s_{\text{pe}}$, $s_{\text{sim}}$ ,$s_{\text{goal}}$ for PPR, BL1 and BL2 trained on temperature sequences (see the supplementary material for more details for this experiment).
Figure 2: Personalization score $s_{\text{pe}}$, similarity score $s_{\text{sim}}$, and goal score $s_{\text{goal}}$ of PPR and baselines BL1 and BL2.
@@ 5.2 Mario Environment
The second setting is the Mario environment provided by the OpenAI gym. This environment replicates the classic “Super Mario Bros” game, where the player-controlled character, Mario, navigates through diverse levels filled with obstacles, enemies, and power-ups and the primary objective is to reach the end of each level. We work with level 1 and for simplicity, in our experiments, Mario can only take two actions for each state: move right or jump right. A player can complete a level through various routes, which results in either winning by reaching the flag or losing by falling off a cliff or touching an enemy (e.g., a Goombas).
@@
Then we construct an original route $\tau_{0}$ where Mario keeps moving right until getting killed by running into a goomba. The original route is generated by always choosing the action of “go right”, thus the route is featured by Mario walking in a straight line on the ground.
@@
To quantitatively assess the extent to which the recourse path aligns with the preferences of the corresponding agent, we calculate the Kullback-Leibler (KL) divergence between the recoursed and original policy, shown in Table 1. To contrast with this “personalized” path, we also compare the policy of the PPR for the coin seeker with that of the coin dodger. In this case, while the generated path recourses the original route, it is not personalized to the coin dodger, thus significantly increasing the KL divergence from 0.342 to 1.049. We do the same for the coin seeker, comparing its policy function with the PPR generated for the coin dodger, and obtain a larger KL divergence of 0.554, up from 0.318 if it was personalized for him.

### v3

@@ 4.1 Reward Shaping
Path-level personalization
Given the edit distance $d(\tau_{0},\tau_{r})$ between two paths $\tau_{0}$ and $\tau_{r}$, as discussed in Section 3.2, the similarity between $\tau_{0}$ and $\tau_{r}$ is computed by
@@
$f_{\text{sim}}(\cdot)$ value is between 0 and 1, where $f_{\text{sim}}(\tau_{0},\tau_{r})=1$ if $\tau_{0}$ and $\tau_{r}$ are identical, and close to 0 if they are very different. $f_{\text{sim}}(\cdot)$ can be used as the path similarity reward $R_{\text{path}}$, which is defined at the sequence level.
Policy-level Personalization
We aim to personalize the path according to the behavior policy $\mathcal{P}_{A}$. This can be achieved by incorporating a personalization reward, $R_{\text{policy}}$, into the reward function. This reward function is derived from agent $A$’s policy $\mathcal{P}_{A}$, which is a probability distribution over state-action pairs. Specifically, for a given state-action pair $(\mathbf{s},a)$, $\mathcal{P}_{A}$ returns a probability representing the likelihood that agent $A$ will choose action $a$ in state $\mathbf{s}$. We utilize this probability to assess whether a recommended action aligns with agent $A$’s behavior, as represented by this distribution.
Directly using the probability as the reward value (i.e., $h(p)=p$) could be misleading because the magnitude of the probabilities is a result of the cardinality of the action space. For example, if $|\mathcal{A}|=2$, a probability of 0.3 is considered low, and the personalization reward should also be low. However, when $|\mathcal{A}|$ is a very large value, 0.3 should be considered large, and the corresponding personalization reward should also be large. Thus, a linear relationship between the reward and probability is insufficient. Therefore, we design a link function to represent the reward
$R_{\text{policy}}(\mathbf{s},a)=h(\mathcal{P}_{A}(\mathbf{s},a)).$ (3)
$h(\cdot)$ needs to satisfy the conditions as follows:
@@
Then, the total policy-level personalization reward defined at the sequence level is the sum of rewards for each state-action pair along the path:
$R_{\text{policy}}(\tau_{r})=\sum_{\mathbf{s}_{t}\in\tau_{r}}R_{\text{policy}}(\mathbf{s}_{t},a_{t}).$ (5)
@@
$R(\tau_{r})=R_{\text{goal}}(\tau_{r})+\lambda_{\text{path}}R_{\text{path}}(\tau_{r},\tau_{0})+\lambda_{\text{policy}}R_{\text{policy}}(\tau_{r}).$ (6)
$\lambda_{\text{path}}$ and $\lambda_{\text{policy}}$ are parameters that determine the weighting of each reward component.
@@ 4.2 Training
We use DQN to train a recourse agent $\mathcal{\pi}_{A}^{r}$ that can generate a recourse path $\tau_{r}$ for $\tau_{0}$ executed by agent $\mathcal{P}_{A}$, such that $\tau_{r}$ maximizes the reward $R(\tau_{r})$. For this paper, we employ the Upper Confidence Bound (UCB) bandit algorithm as our exploration strategy (Auer, Cesa-Bianchi, and Fischer 2002) but PPR is adaptable to different exploration strategies.
@@
We want to encourage the agent to explore states and actions with high state-action values, avoid repetition of the same states/actions, and pay higher priority to the new states/actions. We define an exploration score for taking action $a_{t}$ from a state $\mathbf{s}_{t}$ at time $t$ as follows:
@@
The training algorithm is presented in Algorithm 1 to train $\pi_{A}^{r}$. First, we explore the environment using the strategy in Equation (7) and record all the states, actions as well as corresponding rewards (lines 1-9). Next, we add the $k$ best records to buffer replay $B$ to train network $Q_{\theta}$ and update network $Q_{\theta}^{\prime}$ every $C$ step in the while loop (lines 10-16). Lines 12 converts the reward from trajectory-level to state-action level discounted return to update the network using Equation (1). This process is repeated until it converges, i.e., the agent achieve the optimal reward.
1
Input: Original path $\tau_{0}$, personalized policy $\mathcal{P}_{A}$, an environment parameterized by state set $\mathcal{S}$, action set $\mathcal{A}$, and reward function $R(.)$
@@
2
3Initialize a replay buffer $B$ size $b$
4Initialize two networks $Q_{\theta},Q_{\theta^{\prime}}$ with random parameters $\theta$ and $\theta^{\prime}$
5while not converge do
6       $B=\emptyset$
7      for i=1 to $b$ do
8
9            Sample a path $\tau=<\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{T}>$ using Equation (7) with decay $\epsilon$-greedy strategy
10            Record all the action $a_{t}$ in the path
11            Compute the reward $R(\tau)$ with Equation (6)
@@
13
14      Collect $k$ best paths with the highest $R(\tau)$.
15      foreach $k$ record do
@@ 18      Samples tuples from $B$ and update $\theta$ for $Q_{\theta}$:
($\alpha$ is the learning rate and $L$ is the loss function 1)
19      Set $\theta^{\prime}=\theta$ every $C$ steps in the while loop
@@
Note that PPR is compatible with various other training algorithms, including DQN, A3C, and PPO. In this paper, we use DQN because it performs particularly well in environments with discrete action spaces and efficiently handles situations where interactions with the environment are limited (De La Fuente and Guerra 2024; Henderson et al. 2018).
@@ 4.3 PPR for Supervised Learning
In supervised learning, $\tau$ is not a sequence of states but a sequence data $\mathbf{x}$ with a label $y=Y_{0}$. A personalized recourse means to generate a new sequence $\mathbf{x}^{\prime}$ with a different label $Y_{1}$ (goal), while being similar to $\mathbf{x}$ (path-level) and being in-distribution (policy-level). The goal reward can be defined as the probability that $\mathbf{x}^{\prime}$ is classified as label $Y_{1}$ by a pre-trained classifier. The path-level personalization reward is defined the same way as in Section 4.1, which is the Levenshtein distance. However, other distance metrics such as Euclidean distance and Manhattan distance can also be used if the sequences have fixed lengths. To incorporate path-level personalization reward, we train a generator model from the training data to represent agent $A$. For example, if it is text data, $\mathcal{P}_{A}$ is a language model; if it is a sequence of vectors, $\mathcal{P}_{A}$ can be a simple RNN model that is either pre-trained or provided. Then the personalization policy is defined as the sampling probability of the next symbol, given the current prefix sequence: $\mathcal{P}_{A}(\mathbf{s_{t}},a_{t})=p(\mathbf{u_{t}}|\mathbf{u_{1}},...,\mathbf{u}_{t-1})$, where $\mathbf{u}_{i}$ represents each symbol in sequence data.
@@ 5 Experiments
In this section, we conduct experimental evaluations of PPR in several settings. In each setting, we generate agents with different behavior and then evaluate how PPR generates personalized recourse paths to adapt to those agents. The first two settings are reinforcement learning environments, the third one is text generation, and the last one is supervised learning. Due to the space limit, we included additional experiments and performance evaluations in the supplementary material.
@@
While we do not find any baselines that work for all settings or consider personalization, we find two relevant work that work for some settings but not all, denoted as BL1 (Tsirtsis, De, and Rodriguez 2021) and BL2 (Delaney, Greene, and Keane 2021), which work for some settings but not all. BL1 focuses on generating an alternative sequence of actions differing in at most $k$ actions from the observed sequence that could have potentially a better outcome. It uses dynamic programming to exhaustively search all states and actions. Thus, it is not suitable for applications with very large state and action spaces, such as texts and atari games. BL2, on the other hand, is specifically designed to find counterfactual explanations for labeled sequences with the same lengths. It is used only for classification tasks in supervised learning settings and requires access to the training data and the classifier. Therefore, it is not suitable for reinforcement learning applications or text generation. Both implementation procedures can be found in the supplementary material.
@@
We evaluated the generated paths based on three key aspects: path-level personalization, policy-level personalization, and goal satisfaction, using corresponding scores $s_{\text{policy}},s_{\text{path}},$ and $s_{\text{goal}}$, respectively. The goal satisfaction score $s_{\text{goal}}$ and path score $s_{\text{path}}$ are the same as the goal reward and the path-level reward, i.e., $s_{\text{goal}}=R_{\text{goal}}$ and $s_{\text{path}}=f_{\text{sim}}$ from Equation (2). For policy-level personalization score $s_{\text{policy}}$, we report the average log of probabilities of the generated PPR if it is sampled from the personalized policy $\mathcal{P}_{A}$, which represents the (normalized) likelihood of the new path - the higher this value, the more PPR aligns with the policy, i.e., the more likely the agent will take this path. Thus, for a path $\tau_{r}=<\mathbf{s}_{1}^{r},...,\mathbf{s}_{T}^{r}>$, $s_{\text{policy}}=\frac{1}{T}\sum_{t=1}^{T}\log\>\mathcal{P}_{A}(\mathbf{s}^{r}_{t+1}|\mathbf{s}^{r}_{t},a^{r}_{t})$.
@@ 5.1 Grid-world environment
In this section, we apply PPR to a grid-world environment, shown in Figure 1. In this environment, a taxi driver wants to go to the destination (the white flag) as soon as possible. If the driver reaches the destination, the reward is 80. An extra reward of 30 will be given if the driver can collect money at the dollar sign. The reward for any other step is -1. Originally, the driver picked a bad path with a low reward, which either didn’t collect the money or took a detour to reach the destination. We randomly generate 10 such bad paths and an example is shown in Figure 1. We want to generate a better path that can help the taxi driver get a better reward, which means she will reach the destination faster and collect the money. More importantly, we want to be able to provide personalized recourse path that takes into consideration her driving habit. In addition, the new path should not be too different from the original one to satisfy the path level personalization.
Figure 1: The original path $\tau_{0}$ (the driver didn’t collect reward), the recourse paths for the experienced driver $\tau_{\text{exp}}$ and the recourse path for the new driver $\tau_{\text{new}}$.
We simulate two types of taxi drivers as agents. Driver $A_{\text{exp}}$ represents an experienced driver who is comfortable driving on highways, while driver $A_{\text{new}}$ is a newbie driver who prefers to avoid highways and take local roads instead. We model these two drivers by designing distinct reward functions. For agent $A_{\text{exp}}$, we assign higher rewards for using highway routes and lower rewards for local roads, whereas for agent $A_{\text{new}}$, the rewards are reversed [^3] See the supplementary material for more detailed information.. The policy functions derived from these agents are then provided as inputs to the PPR framework.
We then run Algorithm 1 to generate personalized improved paths for each agent. We set $\lambda_{\text{policy}}=0.1$ and $\lambda_{\text{path}}=0.1$ in Algorithm 1 (see the supplementary material for a sensitivity analysis of $\lambda_{\text{policy}}$ and $\lambda_{\text{path}}$). An example is shown in Figure 1.
As depicted, the recommended path $\tau_{\text{exp}}$, personalized for the experienced driver, passes through the highway area in accordance with the experienced driver’s policy. Similarly, the recommended path $\tau_{\text{new}}$ for the newbie driver follows a local road, aligning with his preferences as reflected by the policy function.
In this example, these new paths are the recourse paths to the original path, as they respect the driver’s driving behavior, facilitate a better outcome for the taxi, and maintain similarity to the original path from the starting point to the destination. Please refer to the supplementary material for all 10 paths and their corresponding recourse paths.
@@
(a) $s_{\text{policy}}$, $s_{\text{path}}$ ,$s_{\text{goal}}$ for PPR and BL1 trained on grid-world
(b) $s_{\text{policy}}$, $s_{\text{path}}$ ,$s_{\text{goal}}$ for PPR, BL1 and BL2 trained on temperature sequences (see the supplementary material for more details for this experiment).
Figure 2: Personalization score $s_{\text{policy}}$, similarity score $s_{\text{path}}$, and goal score $s_{\text{goal}}$ of PPR and baselines BL1 and BL2.
@@ 5.2 Mario Environment
The second setting is the Mario environment provided by the OpenAI gym. This environment replicates the classic “Super Mario Bros” game, where the player-controlled character, Mario, navigates through diverse levels filled with obstacles, enemies, and power-ups and the primary objective is to reach the end of each level. We work with level 1 and for simplicity, in our experiments, Mario can only take two actions for each state: move right or jump right. The game is ended by either winning by reaching the flag or losing by falling off a cliff or touching an enemy (e.g., a Goombas).
@@
Then we construct an original path $\tau_{0}$ where Mario keeps moving right until getting killed by running into a goomba. The original path is generated by always choosing the action of “go right”, thus the path is featured by Mario walking in a straight line on the ground.
@@
To quantitatively assess the extent to which the recourse path aligns with the preferences of the corresponding agent, we calculate the Kullback-Leibler (KL) divergence between the recoursed and original policy, shown in Table 1. To contrast with this “personalized” path, we also compare the policy of the PPR for the coin seeker with that of the coin dodger. In this case, while the generated path recourses the original path, it is not personalized to the coin dodger, thus significantly increasing the KL divergence from 0.342 to 1.049. We do the same for the coin seeker, comparing its policy function with the PPR generated for the coin dodger, and obtain a larger KL divergence of 0.554, up from 0.318 if it was personalized for him.